OpenStudy (anonymous):

help please anyone good in calculus and please explain step by step that would be really great

OpenStudy (ranga):



OpenStudy (ranga):

xy = 600,000 ------ (1) Plot Area A = (x + 60)(y + 100) ----- (2) Put y = 600,000 / x from (1) into (2) Find dA/dx and set it to 0 and solve for x.

OpenStudy (anonymous):

100-36000000/x^2=0 x=600,-600

OpenStudy (anonymous):

is this good

OpenStudy (ranga):

A = (x + 60)(y + 100) = xy + 100x + 60y + 6000 A = 600,000 + 6000 + 100x + 60 * 600,000/x A = 606,000 + 100x + 36,000,000/x dA/dx = 100 - 36,000,000 / x^2 = 0 Yes, x = 600 feet is correct. Find y.

OpenStudy (anonymous):

600000/600=1000=y

OpenStudy (ranga):

Put x = 600 and y = 1000 in (2) and calculate the plot area A.

OpenStudy (anonymous):

(600+60)(1000+100)=726000=A

OpenStudy (ranga):

yes, square feet. Now for the last part of the question, calculate the plot area with the dimensions they have given and find the difference with the minimum plot area.

OpenStudy (anonymous):

(1500+60)(400+100)=780000 780000-726000=54000

OpenStudy (ranga):

There is a little bit of ambiguity in the last question. What if is the other way (1500 + 100)(400+60) = 736,000?

OpenStudy (anonymous):

736000-726000=1000

OpenStudy (ranga):

Yeah, but one cannot be sure how the plant is laid out. We assumed a general x and y and we got x lower than y. But in the last part they just tell you 1500 x 400 without telling the orientation. So if we assume 400 is the x and 1500 is the y we get one answer. If we assume the other way we get a different answer. They should have made it clear in the question. If I had to choose I will choose x = 400 and y = 1500 in keeping with the lower dimension of x we found via calculus.

OpenStudy (anonymous):

yeah the last part is a confusing i think they should put the which one is which but thank u are the best i would not have gotten any where this question if i had to do it alone thank u once again

OpenStudy (ranga):

You are very welcome.